OpenStudy (hang254):

Find the domain of the given function. f(x)=(sqrt (x+8))/((x+4)(x-6))

OpenStudy (hang254):

@lalaly @ParthKohli

OpenStudy (hang254):

@whpalmer4 @goformit100

OpenStudy (hang254):

the answer i got is x ≠ -8, x ≠ -4, x ≠ 6. not sure if its correct though

OpenStudy (goformit100):

It's correct. I suppose

OpenStudy (hang254):

ok, thanks

OpenStudy (whpalmer4):

\[f(x) = \frac{\sqrt{x+8}}{(x+4)(x-6)}\] Why can't \( x = -8\)? \(x=-4\) and \(x=6\) cause a division by 0, so they are restricted values, but what's wrong with \(x=-8\)? What stops you from evaluating the function at that value of \(x\)? Also, what about values of \(x\) where \(x < -8\)? Are they allowed?